four-digit numbers between 3600 and 3699 where one digit equals the sum of the other three
Answer by Edwin McCravy(20063)

The thousands digit is 3, and the hundredths digit is 6.
Let the tens digit be a, and ones digit be b

3+6+a=b or 3+6+b=a or 3+a+b=6 or 6+a+b=3

That is,

9+a=b, 9+b=a, a+b=3, or a+b=-3

The last one is out, so we have

9+a=b, 9+b=a, a+b=3
 
The only solution for 9+a=b is a=0, b=9, which gives 3609

The only solution for 9+b=a is b=0, a=9, which gives 3690

The solution for a+b=3 are

 a=0, b=3, which gives 3603
 a=1, b=2, which gives 3612
 a=2, b=1, which gives 3621
 a=3, b=0, which gives 3630

So all 6 solutions are:

1.  3603
2.  3609
3.  3612
4.  3621
5.  3630
6.  3690

Edwin